OpenStudy (anonymous):

Find the derivative of f'(-1) when f(x)=(x+4/x)^2

OpenStudy (anonymous):

derivative of (blah)^2 = 2(blah) * derivative of blah

OpenStudy (anonymous):

wrong JamesWolf the derivative of f(x)=((x+4)/x)^2 is actually f'(x)= -(8(x+4))/x^3

OpenStudy (anonymous):

james please get off of my question thanks

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

just out of interest, you need to use the chain rule, then the quotient rule. :)

OpenStudy (anonymous):

Lrod, there are rules against just giving the answer, and not showing the method

OpenStudy (anonymous):

yes you have to use the quotient rule

OpenStudy (anonymous):

plus the chain rule, hence my original post derivative of (blah)^2 = 2(blah) * derivative of blah where blah is a friendly term for a function. the quotient rule is needed after that

OpenStudy (anonymous):

well you cant really say blah either :P but yes chain and quo.

OpenStudy (anonymous):

I may have misread and thought you only said 2* the original equation. I am sorry for that

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

I wont use blah again :)

OpenStudy (anonymous):

do you need any help with either of the rules sandy?

OpenStudy (anonymous):

yeah sorry i must have overlooked that portion lol. You still with us Sandy?